OpenStudy (anonymous):

Given that one zero is -4, find all zeros of P(x)=x^3+x^2-22x-40

OpenStudy (anonymous):

factor out the \(x+4\) from your polynomial, then you will have a quadratic, which will probably factor

OpenStudy (anonymous):

Wait where are you getting the x+4?

OpenStudy (anonymous):

in other words, \[xx^3+x^2-22x-40=(x+4)(\text{ some quadratic})\] you can find the quadratic by a) long division (ick) b) synthetic division (very easy) or c) thinking

OpenStudy (anonymous):

if \(-4\) is a zero, then one factor must be \(x+4\)

OpenStudy (anonymous):

\[x+4=0\\ x=-4\] so \(x+4\) is a factor for sure

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

do you know synthetic division?

OpenStudy (anonymous):

we can also do it by thinking if you like

OpenStudy (anonymous):

or we can cheat your choice